OpenStudy (anonymous):

Solve |4x – 10| + 2 ≤ 8

OpenStudy (anonymous):

1≤x≤4

OpenStudy (anonymous):

this is a an inequality right?

OpenStudy (anonymous):

if so its \[1\le x \le4\]

OpenStudy (jdoe0001):

\(\bf |4x - 10| + 2 \le 8\implies |4x - 10| \le 8-2\implies |4x - 10| \le 6\\ \quad \\ |4x - 10| \le 6\implies \begin{cases} +(4x - 10) \le 6\\ \quad \\ \bf -(4x - 10) \le 6 \end{cases}\)